Suppose that the force required to push a saw back and forth 30 times, moving it forward 12cm and back 12cm each time, how much work do you do?
1
Expert's answer
2020-11-02T09:22:19-0500

During the one turn the saw covers the distance "d = 12cm +12 cm = 24cm = 0.24m" moving under the force, say, "F" (in Newtons). Then, the work done in one turn is:


"A_1 = Fd = 0.24F \space [N\cdot m]"

The total work in 30 turn then is:

"A = 30A_1 = 30\times 0.24F \space [N\cdot m] = 7.2F \space [N\cdot m]"

Answer. "7.2F \space [N\cdot m]".